For the vectors shown below, determine the values of A and B that would make A +BF, = 3.5i-4. You may assume that each grid line is 1 unit away from the next. -axis 2 x-axis For the vectors shown below determine the values of A and the acute angle 8 (in degrees that would make the negative y-axis +AT -2.5 -4 Use that g 20 v.-2 and v-3. Note: The angle θ is between the vector and y-axis x-axis 02
Answer #1

v1 = - 7 i - 5 j

v2 = - i + 5 j

now

A v1 + B v2 = - 7 A i - 5 A j - Bi + 5 B j

3.5 i - 4 j = ( - 7 A - B) i + ( 5B - 5A) j

comparing coefficient of i

-7A - B = 3.5... (i)

comparing coefficient of j

-5A + 5B = - 4... (ii)

solving (i) and (ii)

A = - 0.3375

B = - 1.1375

=======

plz post other problem separately

Comment before rate in case any doubt, will reply for sure.. goodluck